# Hamiltonian replica exchange augmented with diffusion-based generative models and importance sampling to assess biomolecular conformational basins and barriers

### v1

@@ Abstract
Enhanced sampling techniques are essential for exploring biomolecular conformational dynamics that occur on timescales inaccessible to conventional molecular dynamics (MD) simulations. This study introduces a framework that combines Hamiltonian replica exchange (REST2) with denoising diffusion probabilistic models (DDPMs) and importance sampling to enhance the mapping of conformational free-energy landscapes. Building on previous applications of DDPMs to temperature replica exchange (TREM), we propose two key improvements. First, we adapt the method to REST2 by treating potential energy as a fluctuating variable. This adaptation allows for more efficient sampling in large biomolecular systems. Second, to further improve resolution in high-barrier regions, we develop an iterative scheme combining replica exchange, DDPM, and importance sampling along known collective variables. Benchmarking on the mini-protein CLN025 demonstrates that DDPM-refined REST2 achieves comparable accuracy to TREM while requiring fewer replicas. Application to the enzyme PTP1B reveals a loop transition pathway consistent with prior complex biased simulations, showcasing the approach’s ability to uncover high-barrier transitions with minimal computational overhead with respect to conventional replica exchange approaches. Overall, this hybrid strategy enables more efficient exploration of free-energy landscapes, expanding the utility of generative models in enhanced sampling simulations.
@@ 1 Introduction
Path-ensemble sampling techniques focus specifically on transition regions10, 11, 12, where standard MD simulations spend only a small fraction of the time. While effective in capturing rare events, these methods typically require prior knowledge of the transition pathway, making initialization nontrivial. Furthermore, reconstructing the underlying kinetics and thermodynamics is challenging, and the generated transition path ensembles tend to be highly correlated, reducing sampling efficiency.
Generalized ensemble sampling provides a third alternative by simulating the system within an altered ensemble where slow transitions are artificially accelerated. This acceleration can be achieved by modifying e.g. the system’s temperature13 or potential energy14. Expanded ensemble techniques such as simulated tempering15 and Gaussian-accelerated MD16 operate with a single system copy, but thermodynamic properties of the unperturbed system are often obtained using replica-based strategies. In these schemes, multiple system replicas experience a ladder of perturbations and periodically exchange configurations, improving sampling and facilitating the reconstruction of thermodynamic properties13, 17, 14.
These methods are, in principle, agnostic to the slow degrees of freedom of a system, which is a significant advantage over biased sampling and even path-ensemble techniques that typically require some prior knowledge of the reaction pathway2. However, this advantage also comes with drawbacks. Because the biasing effort is distributed across many modes of the system, the sampling of each mode is less efficient than in biased sampling, where a single, well-chosen collective variable (CV) is targeted. As a result, large energy barriers cannot be as easily overcome18, 19, making these methods less effective for rare-event sampling when compared to direct biasing along a well-defined CV.
@@
One way to leverage the information contained in higher-temperature replicas—beyond simply providing configurations to the unscaled reference replica—is through reweighting techniques such as T-WHAM20 and MBAR21. More recently, Wang et al. introduced a generative approach based on Denoising Diffusion Probabilistic Models (DDPMs) to refine Temperature Replica Exchange Method (TREM) simulations22. DDPM is a stochastic flow-based method that applies discrete forward and backward diffusion steps between the complex probability distribution in the system’s physical space and a more tractable prior distribution23 in the same space. By learning these noising/denoising processes from imperfect distributions sampled at multiple temperatures, the model can generate new configurations with accurate Boltzmann weights and map them back to the system’s probability distribition22. This approach illustrate a growing trend in the field blending deep-learning methods (including flow-based and diffusion-based generative models) and enhanced sampling molecular simulations24.
@@
Wang et al. demonstrated that DDPM can learn the joint probability distribution in temperature and configuration space from TREM simulations, allowing the generation of new data points with improved accuracy22. Notably, DDPM outperformed MBAR21, particularly in computing free-energy profiles at temperatures beyond the simulated range. More recent score-based approaches25, which refine the diffusion process in a continuous framework, further highlighted the potential of generative models for exploring free-energy landscapes26. A very recent study has also demonstrated their utility (albeit not in a TREM approach) in understanding the conformational free-energy landscapes of intrinsically disordered proteins27.
However, the potential of generative models should not be overstated. These methods do not inherently solve the sampling problem and remain highly sensitive to the probability distribution(s) on which they are trained. If a metastable state is not at least partially explored in the simulation data, the generative model will not be able to discover or sample it24. Consequently, their effectiveness relies heavily on the quality of sampling in the replica exchange simulations.
In that sense, the current key limitations of the DDPM-TREM approach are twofold: first, TREM becomes impractical and computationally expensive for large systems28, 29, and second, large free-energy barriers remain undersampled without additional efforts18, 19. To address these challenges, we propose two key improvements. First, we extend the DDPM framework to the Hamiltonian replica exchange scheme REST214, which scales better with system size and requires less replicas28, 30. We demonstrate that DDPM can learn the joint probability distribution in configuration and rescaled potential energy, performing as effectively as TREM in resolving the conformational free-energy surface. Second, we tackle the sensitivity of DDPM to training data quality, particularly in high-barrier regions corresponding to rare events. For systems where relevant reaction coordinates can be identified, we introduce an iterative scheme that combines replica exchange simulation, DDPM, and importance sampling to progressively refine free-energy surfaces and, in particular, free-energy barriers. We demonstrate the effectiveness of our method with two systems of increasing complexity: the chignolin mini-protein CLN025 as a benchmark test case, and the phosphatase PTP1B. In this last case, we show that our approach can uncover a complex transition pathway between two conformations of an active site loop that has only be characterized very recently using extensive biased approaches31.
@@ 2.1 Basic concepts of DDPM
Denoising Diffusion Probabilistic Models 23, 32 use a stochastic mapping from the empirical complex distribution to a simple prior distribution. DDPMs are inspired by non-equilibrium thermodynamics and are based on two diffusion processes (Figure 1): a forward diffusion process, also called noising, consisting of a fixed Markov chain that progressively adds noise to the data, and a reverse diffusion process, also called denoising, that learns the inverse mapping from noise to the structured data.
Forward diffusion process. Let $\mathbf{x_{0}}$ denote a sample from the original data (consisting of Cartesian coordinates or a set of given collective variables), drawn from the complex probability distribution $q(\mathbf{x_{0}})$. The forward diffusion process progressively adds noise in $T$ diffusion steps, which constitutes a Markov chain made of $T$ states, $\mathbf{x_{t}},t\in\{0,..,T\}$, that become more and more noisy as $t$ increases. The amount of noise added at each step is fixed by a variance schedule $\beta_{1},...,\beta_{T}$. The transition probabilities are Gaussian and are defined as follows:
@@
$L=\mathbb{E}_{q}[D_{KL}(q(x_{T}|x_{0})||p(x_{T}))+\\ \sum_{t>1}D_{KL}(q(x_{t-1}|x_{t},x_{0}||p_{\theta}(x_{t-1}|x_{t}))\\ \ln{p_{\theta}(x_{0}|x_{1})}]\$ (3)
@@ 2.2 Architecture and hyperparameters
We used the same design as presented in Wang et al. 22, based on Ho et al. 23. The deep neural network has a U-Net structure, which is a convolutional neural network consisting of two parts: an encoder (down-sampling path) and a decoder (up-sampling path), with skip connections. We used four down-sampling and four up-sampling residue blocks. The diffusion process consisted of 1000 steps. The network parameters were optimized using the Adam optimizer with a learning rate of $2\times 10^{-5}$, and an exponential moving average with a decay rate of 0.995.
@@ 2.3 Hamiltonian replica exchange with solute tempering REST2
TREM becomes impractical for bigger systems, because the probability of coordinate exchanges between neighboring replicas scales as the square root of the number of degrees of freedom. One way to address this issue is by rescaling the potential energy component of the Hamiltonian, possibly on a fraction of the system (which is of course possible when working with temperature exchange). A popular framework is that of solute tempering, of which the version of Wang et al. 14, called REST2 (Replica Exchange with Solute Tempering 2), is an efficient implementation. In REST2, all replicas are run at the same temperature but on different rescaled potential energy surfaces, rescaling only the solute-solute and solute-water interactions, while keeping the water-water interactions unchanged.
@@ 2.4 Simulation strategy (chignolin)
The system was built and simulated with GROMACS 2019.4 33. We used the NMR structure obtained from the Protein Data Bank (2RVD) 34 of the mutant chignolin CLN025 (YYDPETGTWY). The AMBER99SB-ILDN force field 35 was used, and the system was solvated in a cubic box, containing 1947 TIP3P water molecules and two sodium ions. The minimization was done using the steepest-descent algorithm. A short NVT equilibration (T = 300 K) was performed for 100 ps, followed by a 2-ns NPT equilibration (T = 300 K, P = 1 bar) with a Parrinello-Rahman barostat 36, 37. The final structure was used as the starting structure for all the replicas of the TREM and REST2 simulations. Simulations were performed in the NVT ensemble using the velocity rescaling thermostat with a stochastic term 38. The time step was set to 2 fs. We followed the procedure described in 28: for REST2 simulations, we ran 12 replicas at the reference temperature T = 300 K, with $\lambda$ values yielding temperatures ranging from 290 K to 529 K; for TREM, we ran 32 replicas, at temperatures ranging from 290 K to 600 K. Exchanges were attempted every 5000 timesteps, i.e., every 10 ps.
@@
We used three collective variables: the RMSD of the backbone with respect to the folded structure (taken as the final structure of the NPT equilibration), the radius of gyration computed on the Cα, and the fraction of native contacts $Q$, computed following 39. We used these three collective variables to train a DDPM.
@@
We performed Umbrella Sampling simulations along the RMSD direction, defining 19 windows to ensure good overlap between them. The stiffness constant was set to either 20000 kJ/mol or 40000 kJ/mol depending on the window, and simulations were run for 300 ns for each window. The last 150 ns were considered for analysis. Constraints were applied using Plumed version 2.7.4 40, 41 patched with GROMACS. The free-energy profile along the RMSD was computed using weighted histogram analysis method 42 as implemented in the Grossfield code 43.
@@ 2.5 Simulation strategy (PTP1B)
The system was built from the closed conformation provided by the authors of 31 on their Zenodo repository. We used the AMBER99SB*-ILDN force field 35. The system was solvated with 15298 TIP3P water molecules, 52 sodium ions, and 46 chloride ions, and the energy was minimized using the steepest-descent algorithm. A short NVT equilibration at T = 300 K was performed for 100 ps, followed by a 2-ns NPT equilibration at T = 300 K and P = 1 bar with a Parrinello-Rahman barostat 36, 37. The final structure was used to start all the replicas of the REST2 simulations. The REST2 simulations were performed in the NVT ensemble using the velocity rescaling thermostat with a stochastic term 38, at a reference temperature T = 300 K and in a cubic box of 8 nm. We used 24 replicas with temperature ranging from 290 K to 600 K. Initial REST2 simulations were run for 600 ns with Gromacs 2022.4 44 patched with Plumed version 2.8.1 40, 41. Exchanges were attempted every 5000 timesteps, ie every 10 ps.
@@ 2.6 DDPM-REST2 implementation
We first demonstrate how we can extend the approach of Wang et al. 22 to deal with data from REST2 simulations. In REST2, all replicas are run at the same reference temperature, but on different potential energy surfaces. The potential energy of a configuration $\mathbf{X}$ in replica $i$, is defined as :
@@
In their formulation of DDPM applied to TREM22, Wang et al. proposed to consider temperature as a random fluctuating variable used to learn (and improve) a joint probability distribution in space and temperature. This was trivially achieved using the instantaneous temperature computed from the kinetic energy 22. For REST2, we propose to treat the potential energy as a random fluctuating variable, and to learn the corresponding joint probability distribution in space and potential energy. In practice, we tested two variants in which we either considered either the total potential energy $E_{i}$, or, the rescaled part of it $E_{rescaled,i}(\mathbf{X})=\lambda_{i}E_{pp}(\mathbf{X})+\sqrt{\lambda_{i}}E_{pw}(\mathbf{X})$.
@@
Although DDPM can handle dozens of collective variables22, 26, for large systems some choice need to be made as these cannot encompass e.g. all backbone degrees of freedom. Therefore, even for this small protein for which many more CVs could be used, we trained the DDPM on three CVs deemed important for folding: the RMSD with respect to the folded structure (computed on the backbone atoms), the radius of gyration $R_{g}$ (computed on the $C_{\alpha}$), and the fraction of native contacts $Q$39.
To compute $E_{rescaled,i}$, we need to estimate $E_{pp}$, $E_{pw}$ and $E_{ww}$. For this, we use the rerun command of Gromacs to recalculate the energies along the trajectory of each replica with the different Hamiltonians, which gives for a given structure its potential energy as a function of $\lambda$. We then fit $E_{\lambda_{i}}(\mathbf{X})$ as a function of $\lambda_{i}$ to obtain $E_{pp}$, $E_{pw}$ and $E_{ww}$. Note that other approaches to compute $E_{pp}$, $E_{pw}$ and $E_{ww}$ directly (for example by calculating the energies for groups of atoms containing either only the protein or the solvent) could suffer from issues related to the typical net charge of the protein.
We find that $E_{rescaled,i}$ is Gaussian distributed within a given trajectory, analogous to the kinetic energy (and hence the instantaneous temperature) used in Wang et al. 22, and like the total potential energy function (Figure 2A and B). By computing the minimum free-energy path on the two-dimensional free-energy surface computed along RMSD and $R_{g}$ (see Figure 3 and discussion below), we compare the predictions of DDPM trained with the rescaled part of the potential energy, and the total potential energy (Figure 2C), finding no significant difference (see Methods for details about the training and data generation procedure). We decided to use the former for the rest of this work.

### v2

@@ Abstract
Enhanced sampling techniques are essential for exploring biomolecular conformational dynamics that occur on timescales inaccessible to conventional molecular dynamics (MD) simulations. This study introduces a framework that combines Hamiltonian replica exchange with solute tempering (REST2) with denoising diffusion probabilistic models (DDPMs) and importance sampling to enhance the mapping of conformational free-energy landscapes. Building on previous applications of DDPMs to temperature replica exchange (TREM), we propose two key improvements. First, we adapt the method to REST2 by treating potential energy as a fluctuating variable. This adaptation allows for more efficient sampling in large biomolecular systems. Second, to further improve resolution in high-barrier regions, we develop an iterative scheme combining replica exchange, DDPM, and importance sampling along known collective variables. Benchmarking on the mini-protein CLN025 demonstrates that DDPM-refined REST2 achieves comparable accuracy to TREM while requiring fewer replicas. Application to the enzyme PTP1B reveals a loop transition pathway consistent with prior complex biased simulations, showcasing the approach’s ability to uncover high-barrier transitions with minimal computational overhead with respect to conventional replica exchange approaches. Overall, this hybrid strategy enables more efficient exploration of free-energy landscapes, expanding the utility of generative models in enhanced sampling simulations.
@@ 1 Introduction
Path-ensemble sampling techniques focus specifically on transition regions10, 11, 12, where standard MD simulations spend only a small fraction of the time. While effective in capturing rare events, these methods often require a reasonable initial guess for the transition pathway (and thus even a vague idea about a relevant set of reaction coordinates), making initialization nontrivial. Furthermore, reconstructing the underlying kinetics and thermodynamics is challenging, and the generated transition path ensembles tend to be highly correlated, reducing sampling efficiency. We note however that continuous improvement of these methods has resulted in strategies that are more robust against the above pitfalls, such as replica-exchange transition interface sampling 13, 14.
Generalized ensemble sampling provides a third alternative by simulating the system within an altered ensemble where slow transitions are artificially accelerated. This acceleration can be achieved by modifying e.g. the system’s temperature15 or potential energy16. Expanded ensemble techniques such as simulated tempering17 and Gaussian-accelerated MD18 operate with a single system copy, but thermodynamic properties of the unperturbed system are often obtained using replica-based strategies. In these schemes, multiple replicates of the system experience a ladder of perturbations and periodically exchange configurations, improving sampling and facilitating the reconstruction of thermodynamic properties15, 19, 16.
These methods are, in principle, agnostic to the slow degrees of freedom of a system, which is a significant advantage over biased sampling and even path-ensemble techniques that typically require some prior knowledge of the reaction pathway2. However, this advantage also comes with drawbacks. Because the biasing effort is distributed across many modes of the system, the sampling of each mode is less efficient than in biased sampling, where a single, well-chosen collective variable (CV) is targeted. As a result, large energy barriers cannot be as easily overcome, as we recently illustrated on realistic and complex biomolecular systems20, 21, making these methods less effective for rare-event sampling when compared to direct biasing along a well-defined CV.
@@
One way to leverage the information contained in higher-temperature replicas, beyond simply providing configurations to the unscaled reference replica, is through reweighting techniques such as T-WHAM22 and MBAR23. More recently, Wang et al. introduced a generative approach based on Denoising Diffusion Probabilistic Models (DDPMs) to refine Temperature Replica Exchange Method (TREM) simulations24. DDPM is a stochastic flow-based method that applies discrete forward and backward diffusion steps between the complex probability distribution in the system’s physical space and a more tractable prior distribution25 in the same space. By learning these noising/denoising processes from imperfect distributions sampled at multiple temperatures, the model can generate new configurations with accurate Boltzmann weights and map them back to the system’s probability distribition24. This approach illustrates a growing trend in the field blending deep-learning methods (including flow-based and diffusion-based generative models) and enhanced sampling molecular simulations26.
@@
Wang et al. demonstrated that DDPM can learn the joint probability distribution in temperature and configuration space from TREM simulations, enabling the generation of new data points with improved accuracy24. Notably, DDPM outperformed MBAR23, particularly in computing free-energy profiles at temperatures beyond the simulated range. More recent score-based approaches27, which refine the diffusion process in a continuous framework, further highlight the potential of generative models for exploring free-energy landscapes28. A very recent study has also demonstrated their utility (albeit not in a TREM approach) in understanding the conformational free-energy landscapes of intrinsically disordered proteins29.
However, the potential of generative models should not be overstated. These methods do not inherently solve the sampling problem and remain highly sensitive to the probability distribution(s) on which they are trained. If a metastable state is not at least partially explored in the simulation data, the generative model will not be able to discover or sample it26. Consequently, their effectiveness relies heavily on the quality of sampling in the replica exchange simulations.
In that sense, the current key limitations of the DDPM-TREM approach are twofold: first, TREM becomes impractical and computationally expensive for large systems30, 31, and second, large free-energy barriers remain undersampled without additional efforts20, 21. To address these challenges, we propose two key improvements. First, we extend the DDPM framework to the Hamiltonian replica exchange scheme REST216, which scales better with system size and requires less replicas, and that is regularly used by our group to enhance conformational exploration in biomolecular systems32, 33, 34, 35. We demonstrate that DDPM can learn the joint probability distribution in configuration and rescaled potential energy, performing as effectively as TREM in resolving the conformational free-energy surface. Second, we tackle the sensitivity of DDPM to training data quality, particularly in high-barrier regions corresponding to rare events. For systems where relevant reaction coordinates can be identified, we introduce an iterative scheme that combines replica exchange simulation, DDPM, and importance sampling to progressively refine free-energy surfaces and, in particular, free-energy barriers. We demonstrate the effectiveness of our method with two systems of increasing complexity: the chignolin mini-protein CLN025 as a benchmark test case, and the phosphatase PTP1B. In this last case, we show that our approach can uncover a complex transition pathway between two conformations of an active site loop that has only be characterized very recently using extensive biased approaches36.
@@ 2.1 Basic concepts of DDPM
Denoising Diffusion Probabilistic Models 25, 37 use a stochastic mapping from the empirical complex distribution to a simple prior distribution. DDPMs are inspired by non-equilibrium thermodynamics and are based on two diffusion processes (Figure 1): a forward diffusion process, also called noising, consisting of a fixed Markov chain that progressively adds noise to the data, and a reverse diffusion process, also called denoising, that learns the inverse mapping from noise to the structured data.
Forward diffusion process. Let $\mathbf{x_{0}}$ denote a sample from the original data (consisting of Cartesian coordinates or a set of given collective variables), drawn from the complex probability distribution $q(\mathbf{x_{0}})$. The forward diffusion process progressively adds noise in $T$ diffusion steps, which constitutes a Markov chain made of $T$ states, $\mathbf{x_{t}},t\in\{0,..,T\}$, that becomes more and more noisy as $t$ increases. The amount of noise added at each step is fixed by a variance schedule $\beta_{1},...,\beta_{T}$. The transition probabilities are Gaussian and are defined as follows:
@@
$L=\mathbb{E}_{q}[D_{KL}(q(x_{T}|x_{0})||p(x_{T}))+\\ \sum_{t>1}D_{KL}(q(x_{t-1}|x_{t},x_{0}||p_{\theta}(x_{t-1}|x_{t}))\\ \ln{p_{\theta}(x_{0}|x_{1})}]\\$ (3)
@@
As discussed in ref. 24, the generation process can be conditioned to the temperature of interest $T$.. In practice, we propose a slight change in which we generate points in the CV-space, conditioning them in the temperature-space to reproduce the distribution of "atomistic" temperatures (obtained from the atomistic kinetic energies) seen at $T$ in each replica (which is typically Gaussian-distributed with a variance that scales as $T$). A similar conditioning on the generation process is applied when dealing with the potential energy instead of the kinetic energy in the REST2 framework (see below).
@@ 2.2 Architecture and hyperparameters
We used the same design as presented in Wang et al. 24, based on Ho et al. 25. The deep neural network has a U-Net structure, which is a convolutional neural network consisting of two parts: an encoder (down-sampling path) and a decoder (up-sampling path), with skip connections. We used four down-sampling and four up-sampling residue blocks. The diffusion process consisted of 1000 steps. The network parameters were optimized using the Adam optimizer with a learning rate of $2\times 10^{-5}$, and an exponential moving average with a decay rate of 0.995.
2.3 DDPM error bars
There are several possible ways to define the uncertainties of the DDPM results. First, we trained multiple models on the same datasets and computed the variability among their predictions. In all cases, the resulting errors were very small, at most 0.1 kcal/mol, demonstrating that the intrinsic uncertainties of the model are negligible. The reported error bars instead correspond to the standard deviations obtained from models trained on successive dataset blocks, analogous to the procedure used for estimating raw simulation errors. These deviations are typically comparable to those observed in the underlying data, and therefore reflect the sampling uncertainties of the training data rather than model error per se.
@@ 2.4 Hamiltonian replica exchange with solute tempering REST2
TREM becomes impractical for bigger systems, because the probability of coordinate exchanges between neighboring replicas scales as the square root of the number of degrees of freedom. One way to address this issue is by rescaling the potential energy component of the Hamiltonian, possibly on a fraction of the system (which is of course impossible when working with temperature exchange). A popular framework is that of solute tempering, of which the version of Wang et al. 16, called REST2 (Replica Exchange with Solute Tempering 2), is an efficient implementation. In REST2, all replicas are run at the same temperature but on different rescaled potential energy surfaces, rescaling only the solute-solute and solute-water interactions, while keeping the water-water interactions unchanged.
@@
2.5 Free-energy differences between conformational basins
To compute free-energy differences between basins, we defined boundaries for each basin on the 2D FES and summed the normalized histogram probabilities within each region to the relative probabilities in each state, from which we could finally obtain the free-energy differences.
@@ 2.6 Simulation strategy (chignolin)
The system was built and simulated with GROMACS 2019.4 38. We used the NMR structure obtained from the Protein Data Bank (2RVD) 39 of the mutant chignolin CLN025 (YYDPETGTWY). The Amber99sb-ILDN force field 40 was used, and the system was solvated in a cubic box, containing 1947 TIP3P water molecules and two sodium ions. The minimization was done using the steepest-descent algorithm. A short NVT equilibration (T = 300 K) was performed for 100 ps, followed by a 2-ns NPT equilibration (T = 300 K, P = 1 bar) with a Parrinello-Rahman barostat 41, 42. The final structure was used as the starting structure for all the replicas of the TREM and REST2 simulations. Simulations were performed in the NVT ensemble using the velocity rescaling thermostat with a stochastic term 43. The time step was set to 2 fs. We followed the procedure described in 30: for REST2 simulations, we ran 12 replicas at the reference temperature T = 300 K, with $\lambda$ values yielding temperatures ranging from 290 K to 529 K; for TREM, we ran 32 replicas, at temperatures ranging from 290 K to 600 K. Exchanges were attempted every 5000 timesteps, i.e., every 10 ps. Useful examples and guidelines for running REST2 and TREM simulations with GROMACS can be found in the PLUMED tutorial ecosystem 44.
@@
We used three collective variables: the RMSD of the backbone with respect to the folded structure (taken as the final structure of the NPT equilibration), the radius of gyration computed on the Cα, and the fraction of native contacts $Q$, computed following 45. We used these three collective variables to train a DDPM.
@@
We performed Umbrella Sampling simulations along the RMSD, defining 19 windows to ensure good overlap between them. The stiffness constant was set to either 20000 kJ/mol/nm2 or 40000 kJ/mol/nm2 depending on the window, and simulations were run for 300 ns for each window. The last 150 ns were considered for analysis. Constraints were applied using Plumed version 2.7.4 46, 47 patched with GROMACS. The free-energy profile along the RMSD was computed using weighted histogram analysis method 48 as implemented in the Grossfield code 49.
@@ 2.7 Simulation strategy (PTP1B)
The system was built from the closed conformation provided by the authors of 36 on their Zenodo repository. We used the Amber99sb-ILDNforce field 40. The system was solvated with 15298 TIP3P water molecules, 52 sodium ions, and 46 chloride ions, and the energy was minimized using the steepest-descent algorithm. A short NVT equilibration at T = 300 K was performed for 100 ps, followed by a 2-ns NPT equilibration at T = 300 K and P = 1 bar with a Parrinello-Rahman barostat 41, 42. The final structure was used to start all the replicas of the REST2 simulations. The REST2 simulations were performed in the NVT ensemble using the velocity rescaling thermostat with a stochastic term 43, at a reference temperature T = 300 K and in a cubic box of 8 nm. We used 24 replicas with temperature ranging from 290 K to 600 K. Initial REST2 simulations were run for 600 ns with Gromacs 2022.4 50 patched with Plumed version 2.8.1 46, 47. Exchanges were attempted every 5000 timesteps, ie every 10 ps.
@@ 2.8 DDPM-REST2 implementation
We first demonstrate how we can extend the approach of Wang et al. 24 to deal with data from REST2 simulations. In REST2, all replicas are run at the same reference temperature, but on different potential energy surfaces. The potential energy of a configuration $\mathbf{X}$ in replica $i$, is defined as :
@@
In their formulation of DDPM applied to TREM24, Wang et al. proposed to consider temperature as a random fluctuating variable used to learn (and improve) a joint probability distribution in physical space and temperature. This was trivially achieved using the instantaneous temperature computed from the kinetic energy 24. For REST2, we propose to treat the potential energy as a random fluctuating variable, and to learn the corresponding joint probability distribution in space and potential energy. In practice, we tested two variants in which we considered either the total potential energy $E_{i}$, or, the rescaled part of it $E_{rescaled,i}(\mathbf{X})=\lambda_{i}E_{pp}(\mathbf{X})+\sqrt{\lambda_{i}}E_{pw}(\mathbf{X})$.
@@
Although DDPM can handle dozens of collective variables24, 28, for large systems some choice need to be made as these cannot encompass e.g. all backbone degrees of freedom. Therefore, even for this small protein for which many more CVs could be used, we trained the DDPM on three CVs deemed important for folding: the RMSD with respect to the folded structure (computed on the backbone atoms), the radius of gyration $R_{g}$ (computed on the $C_{\alpha}$), and the fraction of native contacts QQ45. As a consequence, what DDPM generates are points (and thus probability distributions) in this reduced CV-space, and not 3D Cartesian coordinates of the full biomolecule.
To compute $E_{rescaled,i}$, we need to estimate $E_{pp}$, $E_{pw}$ and $E_{ww}$. For this, we use the rerun command of Gromacs to recalculate the energies along the trajectory of each replica with the different Hamiltonians, which gives for a given structure its potential energy as a function of $\lambda$. We then fit $E_{\lambda_{i}}(\mathbf{X})$ as a function of $\lambda_{i}$ to obtain the three unknown variables $E_{pp}$, $E_{pw}$ and $E_{ww}$ for this configuration using Equation 5. Note that other approaches to compute $E_{pp}$, $E_{pw}$ and $E_{ww}$ directly (for example by calculating the energies for groups of atoms containing either only the protein or the solvent) could suffer from issues related to the typical net charge of the protein because of the Ewald summation on a periodic, charged system.
We find that $E_{rescaled,i}$ is Gaussian distributed within a given trajectory, analogous to the kinetic energy (and hence the instantaneous temperature) used in Wang et al. 24, and like the total potential energy function (Figure 2A and B). By computing the minimum free-energy path on the two-dimensional free-energy surface computed along RMSD and $R_{g}$ (see Figure 3 and discussion below), we compare the predictions of DDPM trained with the rescaled part of the potential energy, and the total potential energy (Figure 2C), finding no significant difference (see Methods for details about the training and data generation procedure). We decided to use the former for the rest of this work.